OpenStudy (anonymous):

(p^-4q^10r^4)^5/(p^9q^2r^-8)^-3. Simplify please

OpenStudy (luigi0210):

When you divide exponent powers, you subtract them from one another.

OpenStudy (luigi0210):

So match up variables that are the same and subtract their powers.

OpenStudy (luigi0210):

Also \(\Large (a^b)^c=a^{b*c}\)

OpenStudy (anonymous):

p^47r^44/q^56 is this right? @Luigi0210

OpenStudy (luigi0210):

I got something different

OpenStudy (anonymous):

what did you get?

OpenStudy (luigi0210):

Actually, I'd like to see how you did, then maybe we find out where you went wrong, or if I did.

OpenStudy (anonymous):

Well i am doing this on a online thing and it told me i was wrong so i know its wrong

OpenStudy (luigi0210):

Okay, so here, I'll explain what I did step by step and see if it makes sense to you alright?

OpenStudy (anonymous):

Sounds good!

OpenStudy (anonymous):

are you there? @Luigi0210 \

OpenStudy (luigi0210):

Okay, so first we distribute those exponents outside: \(\LARGE \frac{(p^{-4}q^{10}r^4)^5}{(r^9q^2r^{-8})^{-3}}\) \(\LARGE \frac{p^{-20}q^{50}r^{20}}{p^{-27}q^{-6}r^{24}}\) Make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Now subtract their exponents: \(\Huge p^{-20-(-27)}~q^{50-(-6)}~r^{20-24}\) \(\LARGE p^{7}~q^{56}~r^{-4}\) And I'm here.. OS is just being stupid to me >_>

OpenStudy (anonymous):

lol its fine! so is that the final answer?

OpenStudy (luigi0210):

The simplified answer is \(\LARGE \frac{p^{7}~q^{56}}{r^{4}}\)

OpenStudy (anonymous):

okay thank you so much! i might have more haha

OpenStudy (luigi0210):

You're welcome, and if you have more just open a new question~

OpenStudy (anonymous):

okay!